OpenStudy (bunnyzrockgirl):

slope = 2/5, (10, 3) is on the line. y = mx + b

OpenStudy (bunnyzrockgirl):

@midhun.madhu1987 can you help?

OpenStudy (midhun.madhu1987):

y - y1 = m(x - x1) where x1 = 10 y1 = 3 m=2/5

OpenStudy (bunnyzrockgirl):

thx

OpenStudy (midhun.madhu1987):

Yw :)